If X has a Hypergeometric distribution with sample size (n) = 10, the number of successes in the population (A) = 12, and the population size (N) = 25, then P(X < 3) = ________.
Answer #1

Answer:

Given,

n = 10

A = 12

N = 25

Consider,

P(X = k) = ACk * (N-A)C(n-k) / NCn

P(X < 3) = P(X = 0) + P(X = 1) + P(X = 2)

= 12C0*13C10 / 25C10 + 12C1*13C9 / 25C10 + 12C2*13C8 / 25C10

= 1*286 / 3268760 + 12*715 / 3268760 + 66*1287 / 3268760

= 1/3268760 (286 + 8580 + 84942)

= 96808 / 3268760

= 0.0287